JET-FTIR-SPECTROSCOPY OF PYRROLE CLUSTERS

\underline{INGO DAUSTER}, COREY A. RICE and MARTIN A. SUHM, Institut für Physikalische Chemie, Universität Göttingen, Tammannstr. 6, D-37077 Göttingen, Germany.

Jet-FTIR spectroscopy is a powerful technique to investigate small hydrogen bonded clusters^{a}. The filet-jet-spectrometer combines a 600 × 0.2 mm^{2} slit nozzle with a commercial FTIR-instrument^{b}. Pyrrole is a five-membered heterocyclic aromat with hydrogen bond donor (N–H) and acceptor (π-system) functionalities. Weak N–H⋯ π hydrogen bonds^{c} play an important role in biological systems. Therefore, weakly bound aggregates of pyrrole and some of its methylated derivatives have been systematically characterised by the filet-jet-spectroscopy. In pyrrole-2-carboxaldehyde, the additional carbonyl group offers a better hydrogen bond acceptor site and leads to the formation of more conventional N–H⋯O hydrogen bonds. Its dimer has therefore been investigated as a small model system for hydrogen bonding in peptides.

aT. Häber, U. Schmitt, C. Emmeluth and M. A. Suhm Faraday Discuss. 118, 2001, 331–359.

bC. A. Rice, N. Borho and M. A. Suhm Z. Phys. Chem. 219, 2005, 379–388.

cA. Gómez-Zavagalia and R. Fausto J. Phys. Chem. A 108, 2004, 6953–6967.